OpenStudy (anonymous):

help help.... if y varies jointly as x and z. y=50 when x=4 and z=5 find y when x=8 and z=3

OpenStudy (aroub):

y=kxz 50=k(4)(5) 50=k(20) k=5/2 y=kxz y=5/2(8)(3) y=60